The Windsor Corporation issued 10-year, $5,470,000 par, 7% callable convertible subordinated debentures on January 2, 2017. The bonds have a par value of $1,000, with interest payable annually. The current conversion ratio is 13:1, and in 2 years it will increase to 18:1. At the date of issue, the bonds were sold at 96. Bond discount is amortized on a straight-line basis. Windsor’s effective tax was 35%. Net income in 2017 was $9,350,000, and the company had 1,820,000 shares outstanding during the entire year.

(a) Compute both basic and diluted earnings per share.

Basic earnings per share       $
Diluted earnings per share       $
Answer #1

Answer 1.

basic earning per share = net income / total outstanding shares

basic earning per share = 9,350,000 / 1,820,000 = 5.137 = 5.14

Answer 2.

maturity value = 5470000 * 7% = 382,900 interest

discount amortization = [ ( 1 -.0.96) * 5470000 *1/10] = 21,880

Interest expense = 382900 + 21880  = 404780

1 -tax rate 35% = 65%

404780 * 65% = 263,107

Net income + interest exp = 9,350,000 + 263,107 = 9,613,107

debentures = 5470000 / 1000 = 5470

5470 * 18 = 98460

diluted earning per share = 9613107 / ( 1820000 + 98460) = 5.01